jagr2713 (jagr2713):

Pedro has created the function f(x) = the quantity of 4x-3/2 to represent the number of assignments he has completed, where x represents the number of weeks in the course. Pedro discovers that, using the inverse function to solve for x = 30, he can predict when he will have 30 assignments completed. Explain to Pedro how to accomplish this, using complete sentences.

jagr2713 (jagr2713):

@ganeshie8

jagr2713 (jagr2713):

@KlOwNlOvE

jagr2713 (jagr2713):

@No.name

OpenStudy (anonymous):

30=4x-3/2 cross multiply 60=4x-3 add both sides by 3 63=4x now divide by 4 x=15.75

jagr2713 (jagr2713):

hm thats what i got but idk if it write hee look at the pic http://prntscr.com/458pu4

OpenStudy (anonymous):

Alright I would write Pedro needs to set up the equation like this \[30=\frac{ 4x-3 }{ 2 }\] To solve this he would multiple 30 by 2 to get rid of the fraction. Now the equation should read \[60=4x-3\]. From here he would add 3 into 60 to get rid of the 3. After the equation reads 63=4x, Pedro would need to divide 4 into 63 and get x=15.75. Now from here he would round to the nearest whole number and turn in 30 assignments in 16 days

jagr2713 (jagr2713):

oh ok Thx alot :D

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

thanks for the medal

jagr2713 (jagr2713):

yw